# Purification of Noisy Measurements and Faithful Distillation of Entanglement

### v1

@@ Abstract
Distilling entanglement from pure states. We next consider cases where two parties share pure states which are not maximally entangled. From the Schmidt decomposition, let us suppose that two parties share copies of a state in the following,
$|\psi(\theta)\rangle^{(AB)}=\sin\theta|00\rangle^{(AB)}+\cos\theta|11\rangle^{(AB)}$
where $\theta\in(0,\frac{\pi}{4})$ so that $\cos\theta>\sin\theta$. In a noiseless scenario, Bob implements a local operation with Kraus operators,
@@
If a measurement is noisy, Kraus operators in Eq. (22) cannot be precisely realized and consequently entanglement is not distilled. In particular, two parties may share mixed states due to noise existing in Kraus operators. Similarly to Eq. (23), we have that
@@
where a measurement $\widetilde{M}_{0}^{(E)}$ is noisy, see Eq. (2). Thus, a Kraus operator in Eq. (22) when a measurement is noiseless no longer applies. Even if a measurement outcome 00 occurs, the other rank-one Kraus operator $K_{1}$ appears with a probability $p/2$.
@@
Thus, we have shown distillation of an ebit from a pure state by purifying a noisy measurement. In fact, $F_{n}$ converges to $1$ as $n$ increase, e.g., $p=0.1$ and $\theta=\pi/16$: n 1 2 3 4 $F_{n}$ 0.805 0.984 0.999 1.000 If noise is present in an implementation of a collective CNOT gate, see Eqs. (18) and (19), one cannot reach a unit fidelity. For instance, for a noise fraction $p=0.1$ and shared state $\theta=\pi/16$ we have n 1 2 3 4 $F_{n}$ 0.805 0.914 0.924 0.924 showing that a fidelity converges to $0.924$ with two additional qubits. We detail analytic derivations in Appendix C.
Conclusion. We have shown that a noisy measurement can be purified and applied to the distillation of entanglement. For both cases of mixed and pure states, we have shown that entanglement can be distilled by purifying noisy measurements. The purification protocol is robust against measurement and gate errors in implementation. For a realistic scenario where an error rate is about $5$-$10\%$, we have shown that two additional qubits for the purification suffice almost to recover the noiseless range for distillability: the single fraction of $F>1/2$. Our results are a versatile tool to suppress measurement errors for realizing practical entanglement applications with currently available quantum technologies.
In future investigations, it would be interesting to consider noise in bilateral CNOT gates in the distillation of entanglement. Noise in bilateral CNOT gates prevents two parties from sharing an ebit: even a tiny fraction of noise disallows two parties to establish an ebit. We note that the proposed purification protocol is yet of no use in dealing with errors from bilateral CNOT gates.
This work is supported by the National Research Foundation of Korea (Grant No. NRF-2021R1A2C2006309, NRF-2022M1A3C2069728), the Institute for Information & Communication Technology Promotion (IITP) (RS-2023-00229524), and grant-in-aid of LG Electronics.
Appendix A Derivation of Eq. (16)
In this section, we take different error rates in noisy measurements into account by generalizing Eq.(8), where all measurements have an identical error rate, to cases where each measurement has a different error probability. Let $p_{i}$ denote a noise fraction of the $i$-th measurement, so that
Recall that a POVM element is found as,
In the above, coefficients $r_{0}^{(n)}$ and $r_{1}^{(n)}$ can be obtained recursively from an initial condition $\widetilde{Q}_{x}^{(1)}=\widetilde{M}_{x_{1}}$, where $r_{0}^{(1)}=(1-\frac{p_{1}}{2})$ and $r_{0}^{(1)}=\frac{p_{1}}{2}$, and updating $n$. That is, we exploit the following relation in the update $k=2,\cdots,n$:
where $V_{2}$ is a CNOT gate that takes the first qubit as a control and the $k$-th qubit as a target. From the equations (25, 26), we have
For the generality, let $p_{i}^{R}$ denote a noise fraction of the $i$-th measurement of a party $R\in\{A,B\}$ so that the POVM element above may be written as,
$r_{0}^{A}=\prod_{i=1}^{n}\left(1-\frac{p_{i}^{A}}{2}\right),~{}r_{1}^{A}=\prod_{i=1}^{n}\left(\frac{p_{i}^{A}}{2}\right)~{}~{}\mathrm{and}~{}~{}r_{0}^{B}=\prod_{i=1}^{m}\left(1-\frac{p_{i}^{B}}{2}\right),~{}r_{1}^{B}=\prod_{i=1}^{m}\left(\frac{p_{i}^{B}}{2}\right).$ (27)
$\phantom{=}+\left(\frac{1-F}{3}\right)F\left(r_{\text{even}}(\Pi_{0}-|\phi^{+}\rangle\langle\phi^{+}|)+r_{\text{odd}}\Pi_{1}\right)+\left(\frac{1-F}{3}\right)^{2}\left(r_{\text{odd}}\Pi_{0}+(2r_{\text{even}}+r_{\text{odd}})\Pi_{1}+r_{\text{even}}|\phi^{+}\rangle\langle\phi^{+}|\right),$ (28)
with
$r_{\text{even}}^{(n,m)}$ $=\prod_{i=1}^{n}\left(1-\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(1-\frac{p^{B}_{i}}{2}\right)+\prod_{i=1}^{n}\left(\frac{p^{A}_{i}}{2}\right)\prod_{i=1}^{m}\left(\frac{p^{B}_{i}}{2}\right),~{}\mathrm{and}$ (30)
Appendix B Entanglement distillation with noisy purification protocols: mixed states
In this section, we detail the analysis on the two-way entanglement distillation protocol with the purification of noisy measurements in which CNOT gates are noisy. We show a lower bound of the distillability:
  • i) $F>1/2$ for a noiseless case,
  • ii) $F>L^{(n,m)}$ in Eq. (17) where $L^{(n,m)}$ converges to $1/2$ when $n$ and $m$ tend to be large, and
  • iii) $F>L_{\epsilon}^{(n,m)}$, with $\epsilon$ a noise fraction in a CNOT gate, see Eq. (19), where $L_{\epsilon}^{(n,m)}$ does not converge to $1/2$ for large $n$ and $m$.
To show the item iii), it suffices to consider instances $n=m$. A lower bound found from Eq. (17)
From the relation above, we write by $s$ the limit point,
$s=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}},$
for which it holds that
$s=\frac{(1-\epsilon)s(\frac{p}{2})+\frac{\epsilon}{4}(1+s)}{(1-\epsilon)(1-\frac{p}{2})+\frac{\epsilon}{4}(1+s)}.$
Hence, the limit can be found as
$s$ $=$ $2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}.$
Consequently, when $n$ tends to be large, the lower bound converges as follows,
$\lim_{n\rightarrow\infty}L_{\epsilon}^{(n)}=\frac{2(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(3-2(2-p)p)+\epsilon\sqrt{4(1-p)^{2}(1-2\epsilon)+\epsilon^{2}(5+4(-2+p)p)}}{4(1-\epsilon)^{2}(1-p)^{2}}$
which is strictly away from $1/2$.
@@ Appendix B Entanglement distillation with noisy purification protocols: mixed states
Appendix C Entanglement distillation with noisy purification protocols: pure states
@@
Suppose that we realize the purification of a measurement with $(n-1)$ additional qubits to distill entanglement from a single-copy pure state, where a POVM element is denoted by $\widetilde{Q}_{0}^{(n)}=r_{0}^{(n)}M_{0}+r_{0}^{(n)}M_{1}$. Let $p_{succ}$ denote the probability that the purification of a measurement is successful: we have
@@
If noise fractions $p$ are identical, and CNOT gates in measurement purification have an error $\epsilon$, $r_{0}^{(n)}$ and $r_{1}^{(n)}$ can be calculated by recursive relation:
@@
$s=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}}=2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}$
@@
$\lim_{n\rightarrow\infty}F_{n}$ $=\frac{{2\sin^{2}\theta+\frac{1}{2}s(1-2\sin^{2}\theta)}}{2\sin^{2}\theta+s(1-2\sin^{2}\theta)}$

### v3

@@ Appendix F Distilling entanglement from pure states
We next consider cases where two parties share pure states which are not maximally entangled. From the Schmidt decomposition, let us suppose that two parties share copies of a state, $|\psi(\theta)\rangle^{(AB)}=\sin\theta|00\rangle^{(AB)}+\cos\theta|11\rangle^{(AB)}$ where $\theta\in(0,\frac{\pi}{4})$. In a noiseless scenario, Bob implements a local operation with Kraus operators,
@@
If a measurement is noisy, Kraus operators in Eq. (76) cannot be precisely realized, and consequently, entanglement is not distilled. In particular, two parties may share mixed states due to noise existing in Kraus operators. Similarly to Eq. (77), we have that
@@
where a measurement $\widetilde{M}_{0}^{(E)}$ is noisy. Thus, a Kraus operator in Eq. (76) when a measurement is noiseless no longer applies. Even if a measurement outcome 00 occurs, the other rank-one Kraus operator $K_{1}$ appears with a probability $p/2$.
@@
Thus, we have shown distillation of an ebit from a pure state by purifying a noisy measurement. In fact, $F_{n}$ converges to $1$ as $n$ increases, e.g., $p=0.1$ and $\theta=\pi/16$:
n 1 2 3 4
$F_{n}$ 0.805 0.984 0.999 1.000
@@
Appendix G Distilling entanglement from pure states with noisy purification protocols
@@
Suppose that we realize the purification of a measurement with $(n-1)$ additional qubits to distill entanglement from a single-copy pure state, where a POVM element is denoted by $\widetilde{Q}_{0}^{(n)}=r_{0}^{(n)}M_{0}+r_{1}^{(n)}M_{1}$. Let $p_{succ}$ denote the probability that the purification of a measurement is successful: we have
@@
For the robustness of the purification protocol, we consider noisy CNOT gates in the purification of noisy measurements. A collective CNOT gate $V_{n}$, which may be decomposed as,
$V_{n}^{(S_{1}\cdots S_{n})}=\bigotimes_{j=2}^{n}V_{2}^{(S_{1}S_{j})}$
is factorized with CNOT gates. We consider a noise fraction $\epsilon$ in a CNOT gate
$(1-\epsilon)V_{2}(~{}\cdot~{})V_{2}^{\dagger}+\epsilon\frac{\mathds{1}}{2}\otimes\frac{\mathds{1}}{2}.$
For an identical noise fraction $p$, suppose that CNOT gates in measurement purification have an error $\epsilon$. We obtain recurrence relations for $r_{0}^{(n)}$ and $r_{1}^{(n)}$:
@@
$u=\lim_{n\rightarrow\infty}\frac{r_{1}^{(n)}}{r_{0}^{(n)}}=2(1-p)(1-\frac{1}{\epsilon})+\sqrt{5-4p(2-p)+\frac{4(1-p)^{2}}{\epsilon}(\frac{1}{\epsilon}-2)}$
@@
$\lim_{n\rightarrow\infty}F_{n}$ $=\frac{{2\sin^{2}\theta+\frac{1}{2}u(1-2\sin^{2}\theta)}}{2\sin^{2}\theta+u(1-2\sin^{2}\theta)}$
@@
For instance, for a noise fraction $p=0.1$, $\epsilon=0.05$ and shared state $\theta=\pi/16$ we have
n 1 2 3 4
$F_{n}$ 0.805 0.914 0.924 0.924
showing that a fidelity converges to $0.924$ with two additional qubits.
Appendix H Generalization to high-dimensional systems
The purification protocol can be extended to noisy measurements for high-dimensional quantum systems. Let $\{M_{k}=|k\rangle\langle k|\}_{k=0}^{d-1}$ denote a measurement in the computational basis in a dimension $d$. Suppose that a depolarization noise to a measurement is introduced, for $k\in\{0,\cdots,d-1\}$
$\widetilde{M}_{k}=(1-p)|k\rangle\langle k|+p\frac{\mathds{1}}{d}$
where a depolarization takes all types of noise with equal weights.
The purification protocol can be generalized to $d$-dimensional noisy measurements with a $d$-dimensional XOR gate [40], denoted by $V^{d}$ that works as follows,
$V^{(d)}|a\rangle|b\rangle=|a\rangle|b\oplus a\rangle~{}~{}\mathrm{where}~{}~{}b\oplus a=b+a~{}(\mathrm{mod}~{}d).$
One can also write the gate in the following form,
$V^{(d)}=\sum_{j=0}^{d-1}|j\rangle\langle j|\otimes X_{d}^{j}$
where $X_{d}=\sum_{k=0}^{d-1}|k+1\rangle\langle k|$ is a bit-translation operator in a dimension $d$, referred to as a discrete Weyl operator. Note that $X_{d}^{0}=\mathds{1}$ and $X_{d}^{j}=\sum_{k}|k+j\rangle\langle k|$. For our purpose, we introduce a collective XOR gate on $n$ systems,
$V_{n}^{(d)}=\sum_{j=0}^{d-1}|j\rangle\langle j|\otimes(X_{d}^{j})^{\otimes n-1}.$
The protocol for purifying a noisy measurement $\{\widetilde{M}_{k}\}_{k=0}^{d-1}$ works as follows. For a $d$-dimensional state $\rho$, we prepare $n-1$ $d$-dimensional ancillary systems in a fiducial state $|0\rangle$,
$V_{n}^{(d)}\rho\otimes|0\rangle\langle 0|^{\otimes(n-1)}V_{n}^{(d)\dagger}.$
Noisy measurements are performed on $n$ ancillary qubits. Outcomes are accepted only when $n$ results are identical. Once all outcomes are given by $k$, an effective POVM element $Q_{k}^{(n)}$ can be identified by
$Q_{k}^{(n)}=\frac{(1-(1-\frac{1}{d})p)^{n}}{p_{accept}}M_{k}+\frac{(\frac{p}{d})^{n}}{p_{accept}}\sum_{l\neq k}M_{l}$ (79)
where
$p_{accept}=\sum_{k=0}^{d-1}p(k^{n})=(1-(1-\frac{1}{d})p)^{n}+(d-1)(\frac{p}{d})^{n}.$
Therefore, we have
$Q_{k}^{(n)}\rightarrow M_{k}~{}~{}\mathrm{as}~{}~{}n\rightarrow\infty.$
We have generalized the protocol for purifying a noisy measurement to dimension $d$.